# Improving Nonpreemptive Multiserver Job Scheduling with Quickswap

### v1

@@ 2.3 Polling Systems and Most Servers First
The literature on polling systems is vast [6]. The single-polling-station and infinite-polling-station systems are well-understood [7, 17], and approximations for waiting time in the multiple-polling-station system have been established [8]. Stability issues caused by switchover times have also been studied [18, 16]. However, our multiserver job system considers a mix of single-server and multiserver operations not found in the polling literature. Furthermore, while MSF essentially uses an exhaustive service discipline for switching phases, the class of MSFQ policies uses a more generalized, threshold-based version of exhaustive service that is not analyzed in the prior work. Hence, the analysis of MSFQ in this paper also serves as a new contribution to the literature on polling systems.
@@ 3.1 Multiserver Jobs
We consider a system equipped with $k$ servers. A multiserver job can be represented by an ordered pair $(x,s)$, where $x\in\{1,2,\cdots,k\}$ is the number of servers it needs in order to run and $s$ is the service duration (also known as job size), the time it needs to run on the servers before completion. Jobs occupy a fixed number of servers throughout their time in service, and cannot be preempted: once started, a job must be run until it is complete. We also refer to a job’s server needs as its class. A class-$i$ job is a job that requires $i$ servers. We refer to this job model as the Multiserver Job (MSJ) model.
We consider a stream of multiserver jobs arriving into our system; class-$i$ jobs arrive according to an independent Poisson process with rate $\lambda_{i}$. We assume the service durations of class-$i$ jobs are i.i.d. random variables with exponential distribution, so that $S_{i}\sim\exp(\mu_{i})$.
@@
We define a feasible schedule as a multiset of classes of multiserver jobs that can run in parallel, obeying the rule that the total number of servers requested does not exceed $k$. We use ${\bf u}=(u_{1},u_{2},\cdots,u_{k})$ to denote a feasible schedule in the multiresource system, where it puts $u_{i}$ class-$i$ jobs in service and $\sum\nolimits_{i=1}^{k}i\;u_{i}\leq k,$ as the total server demand cannot exceed $k$.
In the one-or-all setting, $\lambda_{i}=0$ for all $1<i<k$ because jobs can only request one core or all cores in the system. In this case, $\lambda=\lambda_{1}+\lambda_{k}$. A feasible schedule in this case can either be $u_{k}=1$ and $u_{i}=0$ for $i<k$, or $u_{1}\leq k$ and $u_{i}=0$ for $i>1$.
@@ 3.3 General Notation
We define the notation $X\square Y$ as the independent sum of $X$ copies of the random variable $Y$: $X\square Y:=\sum\nolimits_{i=1}^{X}Y_{i},$ where $X\geq 0$ is an integer-valued random variable and the $Y_{i}$’s are i.i.d samples of $Y$. We define the notation $(x)^{+}$ as the positive part of $x$. Specifically, $(x)^{+}:=\max(x,0)$. We define the notation $\widehat{X}(z)$ as the z-transform of the probability mass function of an integer-valued random variable $X$ and $\widetilde{Y}(s)$ as the Laplace-Stieltjes transform of the probability density function of a continuous random variable $Y$.
@@ 4 Policies
In this section, we define the policies we are using in the paper. We first define the Most Servers First (MSF) policy [5, 28], which is known to have shortcomings as discussed in Section 1. Then, based on the MSF policy, we develop the MSF Quickswap (MSFQ) policy for the setting where there are only class-$1$ and class-$k$ jobs in the system. We analyze the mean response time of our MSFQ policy in Section 5.3 and show response time performance improvements with analysis and simulations in Section 6. Inspired by the improvements obtained with the MSFQ policy, we develop the Static Quickswap policy and the Adaptive Quickswap policy, each of which is a generalization of MSFQ to support general job class distributions. We show that the mean response time performance of Static Quickswap and Adaptive Quickswap each compares favorably to MSF in simulations based on real datacenter traces in Section 6.
@@ 4.1 Most Servers First
We analyze the Most Servers First (MSF) policy as described in [5, 28]. Specifically, we define MSF as a non-preemptive policy that favors jobs with the highest server demands. Whenever a job arrives or completes, MSF tries to put as many additional jobs as possible into service, starting with the job that demands the most servers and moving in descending order of server demands. This process ends when either all servers are utilized or MSF has considered all jobs in the queue.
In the one-or-all case where jobs either require 1 server of $k$ servers, MSF has a somewhat simpler structure. At any moment in time, the policy either serves 1 class-$k$ job or up to $k$ class-$1$ jobs. The two job classes are never served simultaneously. We describe this structure by saying that MSF undergoes two phases. In phase 1, MSF serves exclusively class-$k$ jobs one at a time. Doing so, MSF uses all servers in the system and is thus very efficient. In phase 2, MSF serves exclusively class-$1$ jobs. During phase 2, there can be up to $k$ class-$1$ jobs in service, depending on how many class-$1$ jobs are in the system. Whenever the number of class-1 jobs in service is less than $k$, some of the servers’ service capacity is wasted. MSF only switches between phases when it runs out of jobs of the current class.
@@ 4.2 Most Servers First Quickswap (MSFQ)
To reduce the load-amplifying effect found in MSF, we introduce the Most Server First Quickswap (MSFQ) policy in the one-or-all setting. The goal of MSFQ is to shorten the duration of phase 2 of the MSF policy so that fewer class-$k$ jobs are allowed to build up during this phase. Specifically, an MSFQ policy is associated with a threshold, $\ell$, that is used to shorten the periods where class-$1$ jobs are in service. When the number of class-1 jobs drops below $\ell$, the system stops allowing class-1 arrivals to enter service.
We define the MSFQ policy formally via the following phases with a threshold $\ell\in[0,n-1]$:
@@
At the end of phase 4, the server returns to phase 1.
By analyzing response time of the MSFQ system in Section 5.3, we find that the mean response time is dependent on the number of jobs in the system at the beginning of the phase, and the phase duration — the amount of time from when the system enters phase $i$ until phase $i$ completes. We let random variable $N_{i}^{S}$ denote the number of small (class-1) jobs and $N_{i}^{L}$ denote the number of large (class-$k$) jobs at the beginning of phase $i$. We also let the random variable $H_{i}$ denote the duration of the $i$th phase.
Note that, when $\ell=0$, our MSFQ policy is the same as the MSF policy. We show that the class of MSFQ policies, irrespective of the threshold $\ell$, is throughput optimal in Section 5.1. The intuition behind throughput-optimality follows from the fact that we always fully utilize all the cores outside of the switchover period (phases 3 and 4) when there are enough jobs waiting in the system. By searching over the choices of $\ell$, we show that our MSFQ policy offers significant performance improvement over MSF. This confirms that the load-amplifying effect found in MSF has a huge impact on response time performance.
@@ 4.3 Static Quickswap
Our one-or-all MSFQ analysis relies heavily on the fact that there are only two classes of jobs in the system. The key difficulty in generalizing the MSFQ policy to general job class distributions lies in selecting the next job to serve, once the phase corresponding to the current class of jobs is complete.
We therefore define the Static Quickswap scheduling policy, which cycles through all classes of jobs in a fixed order, with the following two phases for each class of jobs:
  • Working Phase: Run schedule $u_{j}=\lfloor k/i\rfloor\mathbbm{1}\{j=i\}$ and serve class-$i$ jobs exclusively until the number of idle servers exceeds $k$-$\ell$.
  • Draining Phase: Complete the class-$i$ jobs that are already in service. New class-$i$ arrivals are not allowed to enter service during this phase.
@@ 5 Analysis of the MSFQ Policy
In this section, we analyze the Most Servers First Quickswap (MSFQ) policy under the one-or-all setting, where there are only class-1 and class-$k$ jobs. We call the class-1 jobs small jobs and the class-$k$ jobs large jobs. We begin by proving that the MSFQ policy is throughput-optimal in the simplified setting.
@@
[MSFQ Throughput-Optimality]theoremmsfqstability In the 2-class one-or-all MSJ system, the Most Servers First with Quickswap policy has optimal stability region for all thresholds $\ell$, $0\leq\ell<n$.
We then analyze the mean response time under MSFQ. We observe that the mean response time depends on two central factors: the amount of time the policy spends in each phase, and how many jobs are in the system at the beginning of each phase. We therefore approximate the Laplace-Stieltjes transforms of the distributions of phase durations and the z-transforms of the number of jobs distributions at the beginning of each phase. We then show how to use these transforms to approximate the mean response time for small and large jobs. These three steps give the following summary theorem regarding $\mathbb{E}[T]$.
@@
[MSFQ Summary Theorem]theoremmsfqsummary The mean response time under MSFQ, $\mathbb{E}[T]$, depends on the first and second moments of $H_{i}$ and $N_{i}$ for all phases $i$. Hence, one can compute $\mathbb{E}[T]$ using the transforms $\widehat{N_{i}^{S}}(z)$ and $\widehat{N_{i}^{L}}(z)$ for all $i$ using Lemmas 5.3-5.3.
@@ 5.1 Throughput-optimality of MSFQ
We now prove that the MSFQ policy achieves optimal stability region, for any Quickswap threshold parameter $\ell$. We start by lower-bounding its stability region:
@@
In the 2-class one-or-all MSJ system, the Most Servers First with Quickswap policy is positive recurrent for all thresholds $\ell$, $0\leq\ell<k$, whenever $\frac{\lambda_{1}}{k\mu_{1}}+\frac{\lambda_{k}}{\mu_{k}}<1.$

### v2

@@ 2.3 Polling Systems and Most Servers First
The literature on polling systems is vast [borst2018polling]. The single-polling-station and infinite-polling-station systems are well-understood [borst1997polling, foss1996polling], and approximations for waiting time in the multiple-polling-station system have been established [borst1998waiting]. Stability issues caused by switchover times have also been studied [fricker1994monotonicity, foss1996dominance]. However, our multiserver job system considers a mix of single-server and multiserver operations not found in the polling literature. Furthermore, while MSF essentially uses an exhaustive service discipline for switching phases, the class of MSFQ policies uses a more generalized, threshold-based version of exhaustive service that is not analyzed in the prior work. Hence, the analysis of MSFQ in this paper also serves as a new contribution to the literature on polling systems.
@@ 3.1 Multiserver Jobs
We consider a system with $k$ servers. A multiserver job can be represented by an ordered pair $(i,s)$, where $i\in\{1,2,\cdots,k\}$ is the number of servers the job needs in order to run and $s$ is the service duration (also known as job size), the time the job must run on the servers before completion. Jobs occupy a fixed number of servers throughout their time in service, and cannot be preempted: once started, a job must be run until it is complete. We refer to this job model as the Multiserver Job (MSJ) model.
The MSJ model reflects the realities of scheduling in modern large-scale compute clusters. Specifically, running jobs typically cannot be preempted because they are stateful and preemption would destroy this working state [psychas2017non]. Additionally, the MSJ model does not aim to capture straggler effects where a job’s tasks on some servers finish earlier than others. While the straggler effect is captured by more detailed models such as fork-join queueing models [forkjoin], these models are notoriously intractable to analyze. Furthermore, modern systems employ a variety of techniques to mitigate the straggler effects [ananthanarayanan2013effective]. As a result, real-world systems such as Google Borg make scheduling decisions based on fixed server needs and largely ignore straggler effects within a job.
We consider workloads composed of different job classes, where class-$i$ jobs all need $i$ servers. We consider serving a stream of multiserver jobs, where class-$i$ jobs arrive according to an independent Poisson process with rate $\lambda_{i}$. We further assume the service durations of class-$i$ jobs are i.i.d. exponentially distributed random variables such that $S_{i}\sim\exp(\mu_{i})$ for any class $i$.
@@
We define a feasible schedule as a multiset of classes of multiserver jobs that can run in parallel, obeying the rule that the total number of servers requested does not exceed $k$. We use ${\bf u}=(u_{1},u_{2},\cdots,u_{k})$ to denote a feasible schedule in the multiserver system, where it puts $u_{i}$ class-$i$ jobs in service and $\sum\nolimits_{i=1}^{k}i\;u_{i}\leq k,$ as the total server demand cannot exceed $k$.
In the one-or-all setting, $\lambda_{i}=0$ for all $1<i<k$ because jobs can only request one server or all servers in the system. In this case, $\lambda=\lambda_{1}+\lambda_{k}$. A feasible schedule in this case can either be $u_{k}=1$ and $u_{i}=0$ for $i<k$, or $u_{1}\leq k$ and $u_{i}=0$ for $i>1$.
@@ 3.3 General Notation
We define the notation $\Sigma(X,Y)$ as the independent sum of $X$ copies of the random variable $Y$: $\Sigma(X,Y):=\sum\nolimits_{i=1}^{X}Y_{i},$ where $X\geq 0$ is an integer-valued random variable and the $Y_{i}$’s are i.i.d samples of $Y$. We define the notation $(x)^{+}$ as the positive part of $x$. Specifically, $(x)^{+}:=\max(x,0)$. We define the notation $\widehat{X}(z)$ as the z-transform of the probability mass function of an integer-valued random variable $X$ and $\widetilde{Y}(s)$ as the Laplace-Stieltjes transform of the probability density function of a continuous random variable $Y$.
@@ 4 Policies
In this section, we define the policies we are using in the paper. We first define the Most Servers First (MSF) policy [beloglazov, maguluri-2012], which is known to have shortcomings as discussed in Section 1. Then, based on the MSF policy, we develop the MSF Quickswap (MSFQ) policies for the setting where there are only class-$1$ and class-$k$ jobs in the system. We analyze the mean response time of an MSFQ policy in Section 5.3 and show response time performance improvements with analysis and simulations in Section 6. Inspired by the improvements obtained with the MSFQ policies, we develop the Static Quickswap policy and the Adaptive Quickswap policy, each of which is a generalization of MSFQ to support arbitrary sets of job classes. We show that the mean response time performance of Static Quickswap and Adaptive Quickswap each compares favorably to MSF in simulations based on real datacenter traces in Section 6.
@@ 4.1 Most Servers First
We analyze the Most Servers First (MSF) policy as described in [beloglazov, maguluri-2012]. Specifically, we define MSF as a non-preemptive policy that favors jobs with the highest server demands. Whenever a job arrives or completes, MSF tries to put as many additional jobs as possible into service, starting with the job that demands the most servers and moving in descending order of server demands. This process ends when either all servers are utilized or MSF has considered all jobs in the queue.
In the one-or-all case where jobs either require 1 server or $k$ servers, MSF has a somewhat simpler structure. At any moment in time, the policy either serves 1 class-$k$ job or up to $k$ class-$1$ jobs. The two job classes are never served simultaneously. We describe this structure by saying that MSF undergoes two phases. In phase 1, MSF serves exclusively class-$k$ jobs one at a time. Doing so, MSF uses all servers in the system and is thus very efficient. In phase 2, MSF serves exclusively class-$1$ jobs. During phase 2, there can be up to $k$ class-$1$ jobs in service, depending on how many class-$1$ jobs are in the system. Whenever the number of class-1 jobs in service is less than $k$, some of the servers’ service capacity is wasted. MSF only switches between phases when it runs out of jobs of the current class.
@@ 4.2 Most Servers First Quickswap (MSFQ)
To reduce the load-amplifying effect found in MSF, we introduce the Most Server First Quickswap (MSFQ) policies in the one-or-all setting. The goal of MSFQ is to shorten the duration of phase 2 of the MSF policy so that fewer class-$k$ jobs are allowed to build up during this phase. Specifically, an MSFQ policy is associated with a threshold, $\ell$, that is used to shorten the periods where class-$1$ jobs are in service. When the number of class-1 jobs drops below $\ell$, the system stops allowing class-1 arrivals to enter service.
We define the MSFQ policy formally via the following phases with a threshold $\ell\in[0,k-1]$:
@@
At the end of phase 4, the server returns to phase 1. Note that phase 2 and phase 3 are similar, but it will be convenient to treat them separately in our analysis.
By analyzing response time of the MSFQ system in Section 5.3, we find that the mean response time is dependent on the number of jobs in the system at the beginning of the phase, and the phase duration — the amount of time from when the system enters phase $i$ until phase $i$ completes. We let random variable $N_{i}^{L}$ denote the number of light (class-1) jobs and $N_{i}^{H}$ denote the number of heavy (class-$k$) jobs at the beginning of phase $i$. We also let the random variable $H_{i}$ denote the duration of the $i$th phase.
Note that, when $\ell=0$, our MSFQ policy is the same as the MSF policy. We show that the class of MSFQ policies, irrespective of the threshold $\ell$, is throughput optimal in Section 5.1. The intuition behind throughput-optimality follows from the fact that we always fully utilize all the servers outside of the switchover period (phases 3 and 4) when there are enough jobs waiting in the system. By searching over the choices of $\ell$, we show that our MSFQ policy offers significant performance improvement over MSF. This confirms that the load-amplifying effect found in MSF has a huge impact on response time performance.
@@ 4.3 Static Quickswap
Our one-or-all MSFQ analysis relies heavily on the fact that there are only two classes of jobs in the system. The key difficulty in generalizing the MSFQ policies to arbitrary sets of job classes lies in selecting the next job to serve, once the phase corresponding to the current class of jobs is complete.
We therefore define the Static Quickswap scheduling policy, which cycles through all classes of jobs in a fixed order, with the following two phases for each class of jobs, $i$:
  • Working Phase: In class $i$’s working phase, we serve class-$i$ jobs exclusively until the number of idle servers exceeds $k$-$\ell$. Formally, in class $i$’s working phase we set $u_{i}=\lfloor k/i\rfloor$, and we set $u_{j}=0$ for all $j\neq i$.
  • Draining Phase: During class $i$’s draining phase, we complete the class-$i$ jobs that are still in service at the end of the working phase. New class-$i$ arrivals are not allowed to enter service during this phase.
@@ 5 Analysis of the MSFQ Policy
In this section, we analyze the Most Servers First Quickswap (MSFQ) policies under the one-or-all setting, where there are only class-1 and class-$k$ jobs. We call the class-1 jobs light jobs and the class-$k$ jobs heavy jobs. We begin by proving that any MSFQ policy is throughput-optimal in the simplified setting.
@@
[MSFQ Throughput-Optimality]theoremmsfqstability In the one-or-all MSJ system, Most Servers First with Quickswap (MSFQ) policies have optimal stability region for all thresholds $\ell$, $0\leq\ell<k$.
We then analyze the mean response time under MSFQ. We observe that the mean response time depends on two central factors: the amount of time the policy spends in each phase, and how many jobs are in the system at the beginning of each phase. We therefore approximate the Laplace-Stieltjes transforms of the distributions of phase durations and the z-transforms of the number of jobs distributions at the beginning of each phase. We then show how to use these transforms to approximate the mean response time for light and heavy jobs. These three steps give the following summary theorem regarding $\mathbb{E}[T]$.
@@
[MSFQ Summary Theorem]theoremmsfqsummary The mean response time under MSFQ, $\mathbb{E}[T]$, depends on the first and second moments of $H_{i}$ and $N_{i}$ for all phases $i$. Hence, one can compute $\mathbb{E}[T]$ using the transforms $\widehat{N_{i}^{L}}(z)$ and $\widehat{N_{i}^{H}}(z)$ for all $i$ using Lemmas 5.3-5.3.
@@ 5.1 Throughput-optimality of MSFQ
We now prove that MSFQ policies achieve optimal stability region, for any Quickswap threshold parameter $\ell$. We start by lower-bounding its stability region: {restatable}theoremstability In the one-or-all MSJ system, the Most Servers First with Quickswap policies are positive recurrent for all thresholds $\ell$, $0\leq\ell<k$, whenever $\frac{\lambda_{1}}{k\mu_{1}}+\frac{\lambda_{k}}{\mu_{k}}<1.$
Proof.
We prove this theorem via the Foster-Lyapunov theorem with a carefully designed Lyapunov function. See A for a full proof. ∎
@@
In the one-or-all MSJ system, no scheduling policy is stable if $\lambda_{1}/k\mu_{1}+\lambda_{k}/\mu_{k}\geq 1,$